The angle between the tangent drawn from the point (1,4) to the parabola y^2=4x is?

Grade:12

1 Answers

Sayam
6 years ago
Angle is 60 degrees
As y2=4x, so a=1
the equation of tangent is y=mx+1/m.....(i)
it passes through (1,4).
so the equation (i) becomes m2-4m+1=0
solving the quadratic and substituting its value in formula tan theta=m2-m1 /1+m2m1 , you will get your answer. hope this clears your doubt. thanks